jhonyy9 (jhonyy9):

can be writing every primes greater than 2 in the form 2x+1 - if x is one number from the set of natural numbers greater or equal 1- is it this true ?

OpenStudy (anonymous):

yep, buddy! :)

OpenStudy (anonymous):

No prime is even except 2. So each prime except 2 is of the form \[2n+1\]

jhonyy9 (jhonyy9):

ok so if this is true resulted that if p is number prime p=2x+1 and k is number prime k=2y+1 from this k+p=(2y+1)+(2x+1) such that k+p=2x+2y+1+1 what will be k+p=2(x+y)+2 --- k+p=2(x+y+1) --- is it this true ?

OpenStudy (anonymous):

yup tats also true! :) sum of 2 primes (each or one being not 2) is always even..:)